OpenStudy (anonymous):

solve x^2 + 7 = -3 using quadratic formula

OpenStudy (kirbykirby):

\[ x^2+7 = -3\\ x^2+7+3=0\\ x^2+10 = 0\] If you want, you can re-write this as: \[ x^2+0x+10=0\], so then: \[ x=\frac{-(0)\pm\sqrt{0^2-4(1)(10)}}{2(1)}\\ x=\frac{\pm\sqrt{-40}}{2}\\ x=\frac{\pm i\sqrt{40}}{2}=\frac{\pm2i\sqrt{10}}{2}=\pm i\sqrt{10}\]